OpenStudy (anonymous):

is my answer right? i think its option 3

OpenStudy (anonymous):

How can this multiplication problem be written using only improper fractions? 5* 2 1/8= ? 5/1 * 16/8 =? 5/1 * 8/6= ? 5/1 * 17/8=? 5/1 * 18/=?

OpenStudy (turingtest):

you are right

OpenStudy (anonymous):

oh ok thanks :D